# Multi-Derivative Runge-Kutta Flux Reconstruction for hyperbolic conservation laws

### v2

@@ 1 Introduction
which is obtained after spatial discretization of a PDE, Runge-Kutta methods make use of only the right hand side function $\bm{L}$. Multiderivative Runge-Kutta (MDRK) [19] methods were initially developed to solve systems of ODE that also make use temporal derivatives of $\bm{L}$, see [26] for a review of MDRK methods for solving ODEs. They were first used for temporal discretization of hyperbolic conservation laws in [26] by using Weighted Essentially Non-Oscillatary (WENO) [29] and Discontinuous Galerkin (DG) [9] methods for spatial discretization.
@@ 2 Conservation law and solution space
where $\bm{u}\in\mathbb{R}^{p}$ is the vector of conserved quantities, $\bm{f}(\bm{u})$ is the corresponding flux, together with some initial and boundary conditions. The solution that is physically correct is assumed to belong to an admissible set, denoted by $\mathcal{U}_{\textrm{$\operatorname{ad}$}}$. For example in case of compressible flows, the density and pressure (or internal energy) must remain positive. In case of shallow water equations, the water depth must remain positive. In most of the models that are of interest, the admissible set is a convex subset of $\mathbb{R}^{p}$, and can be written as
@@
$\Omega_{e}=[x_{{e-\frac{1}{2}}},x_{{e+\frac{1}{2}}}]\qquad\operatorname{and}\qquad\Delta x_{e}=x_{{e+\frac{1}{2}}}-x_{{e-\frac{1}{2}}}$
@@
which will be taken to be Gauss-Legendre (GL) or Gauss-Lobatto-Legendre (GLL) nodes, and will also be referred to as solution points. There are associated quadrature weights $w_{j}$ such that the quadrature rule is exact for polynomials of degree up to $12N+$ 1 for GL points and upto degree $12N-$ 1 for GLL points. Note that the nodes and weights we use are with respect to the interval $[0,1]$ whereas they are usually defined for the interval $[-1,+1]$. The solution inside an element is given by
@@ 3 Runge-Kutta FR
We first give a brief resume of RKFR scheme since our method has some similarities and borrows ideas from FR method. The RKFR scheme is based on an FR spatial discretization leading to a system of ODE followed by application of an RK scheme to march forward in time. The key idea is to construct a continuous polynomial approximation of the flux which is then used in a collocation scheme to update the nodal solution values. At some time $t$, we have the piecewise polynomial solution defined inside each cell; the FR scheme can be described by the following steps.
@@
$\bm{f}_{h}(\xi,t)=\left[\bm{f}_{{e-\frac{1}{2}}}(t)-\bm{f}_{h}^{\delta}(0,t)\right]g_{L}(\xi)+\bm{f}_{h}^{\delta}(\xi,t)+\left[\bm{f}_{{e+\frac{1}{2}}}(t)-\bm{f}_{h}^{\delta}(1,t)\right]g_{R}(\xi)$
@@
We obtain the system of ODE by collocating the PDE at the solution points
$\frac{\text{\text{{{d}}}}\bm{u}_{p}^{e}}{\text{\text{{{d}}}}t}(t)=-\frac{1}{\Delta x_{e}}\frac{\partial\bm{f}_{h}}{\partial\xi}(\xi_{p},t),\qquad 0\leq p\leq N$
@@ 4 Multi-derivative Runge-Kutta FR scheme
Multiderivative Runge-Kutta [19] methods were initially developed to solve systems of ODE like
@@
that use temporal derivatives of $\bm{L}$. In this work, we use the two stage fourth order multiderivative Runge-Kutta method from [18]. For the system of ODE (4), the MDRK scheme of [18] to evolve from $t^{n}$ to $t^{n+1}$ is given by
@@
$\begin{split}\bm{F}&:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)\\ \bm{F}^{\ast}&:=\bm{f}(\bm{u}^{n})+\frac{1}{6}\Delta t\left(\frac{\partial}{\partial t}\bm{f}(\bm{u}^{n})+2\frac{\partial}{\partial t}\bm{f}(\bm{u}^{\ast})\right)\end{split}$ (7)
@@ 4.1 Conservation property
The computation of correct weak solutions for non-linear conservation laws in the presence of discontinuous solutions requires the use of conservative numerical schemes. The Lax-Wendroff theorem shows that if a consistent, conservative method converges, then the limit is a weak solution. The method (8) is also conservative though it is not directly apparent; to see this multiply (8) by the quadrature weights associated with the solution points and sum over all the points in the $e^{\operatorname{th}}$ element,
$\begin{split}\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{\ast}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{2\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}_{h}}{\partial\xi}(\xi_{p})\\ \sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n+1}&=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\sum_{p=0}^{N}w_{p}\frac{\partial\bm{F}^{\ast}_{h}}{\partial\xi}(\xi_{p})\end{split}$
@@ 4.3 Approximate Lax-Wendroff procedure
$\bm{F}:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t\frac{\partial}{\partial t}\bm{f}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\left(\bm{u}\right)\text{\text{{{d}}}}t$ (12)
@@ 5.1 Stage 1
$\mathsf{\bm{F}}_{e}=a\mathsf{\bm{U}}_{e}\quad\operatorname{where}\quad\mathsf{\bm{U}}_{e}=\mathsf{T}^{(1)}\mathsf{u}_{e}\quad\operatorname{and}\quad\mathsf{T}^{(1)}=\mathsf{I}-\frac{\sigma}{4}\mathsf{D}$
@@ 5.2 Stage 2
the flux derivative at the solution points is given by
@@ 6 Blending scheme
The MDRK scheme (8) gives a high (fourth) order method for smooth problems, but most practical problems involving hyperbolic conservation laws consist of non-smooth solutions like shocks. In such situations, using a higher order method is bound to produce Gibbs oscillations, as stated in Godunov’s order barrier theorem [14]. The cure is to non-linearly add dissipation in regions where the solution is non-smooth, with methods like artificial viscosity, limiters and switching to a robust lower order scheme; the resultant scheme will be non-linear even for linear equations. In this work, we develop a subcell based blending scheme for MDRK similar to the one in [4] for the high order single stage Lax-Wendroff methods. The limiter is applied to each MDRK stage.
@@
where $\alpha_{e}\in[0,1]$ must be chosen based on the local smoothness indicator. If $\alpha_{e}=0$ then we obtain the high order MDRK scheme, while if $\alpha_{e}=1$ then the scheme becomes the low order scheme that is less oscillatory. In subsequent sections, we explain the details of the lower order scheme and the design of smoothness indicators. The lower order scheme will either be a first order finite volume scheme or a high resolution scheme based on MUSCL-Hancock idea. In either case, the common structure of the low order scheme can be explained as follows. In the first stage, the lower order residual $\mathsf{R}^{\ast,L}_{e}$ performs evolution from time $t^{n}$ to $t^{n+\frac{1}{2}}$ while, in the second stage, $\mathsf{R}^{L}_{e}$ performs evolution from $t^{n}$ to $t^{n+1}$. It may seem more natural to evolve from $t^{n+\frac{1}{2}}$ to $t^{n+1}$ in the next stage, but that will violate the conservation property because of the expression of second stage of MDRK (625).
@@
The MDRK scheme used in this work is fourth order accurate so we use polynomial degree $N=3$ in what follows, but write the procedure for general $N$. Let us subdivide each element $\Omega_{e}$ into $N+1$ subcells associated to the solution points $\{x^{e}_{p},p=0,1,\ldots,N\}$ of the MDRK scheme. Thus, we will have $N+2$ subfaces denoted by $\{x^{e}_{p+\frac{1}{2}},p=-1,0,\ldots,N\}$ with $x^{e}_{-\frac{1}{2}}=x_{{e-\frac{1}{2}}}$ and $x^{e}_{N+\frac{1}{2}}=x_{{e+\frac{1}{2}}}$. For maintaining a conservative scheme, the $p^{\operatorname{th}}$ subcell is chosen so that
@@
where $w_{p}$ is the $p^{\operatorname{th}}$ quadrature weight associated with the solution points. Figure 2 gives an illustration of the subcells for degree $N=3$ case. We explain the blending scheme for the second stage, where the lower order scheme performs evolution from $t^{n}$ to $t^{n+1}$. The low order scheme is obtained by updating the solution in each of the subcells by a finite volume scheme,
@@ 6.3 Flux limiter
The flux limiter to enforce admissibility will be applied in each stage so that the evolution of each stage is admissibility preserving. Thus, the admissibility enforcing flux limiting procedure of [4] applied to each stage of MDRK ensures admissibility of that stage and thus of the overall scheme. Since the blending is performed at each stage, the procedure remains the same as in the single stage method of [4] and the reader is referred to the same for details.

### v3

@@ 1 Introduction
which is obtained after spatial discretization of a PDE, Runge-Kutta methods make use of only the right hand side function $\bm{L}$. Multiderivative Runge-Kutta (MDRK) [22] methods were initially developed to solve systems of ODEs that also make use temporal derivatives of $\bm{L}$, see [31] for a review of MDRK methods for solving ODEs. They were first used for temporal discretization of hyperbolic conservation laws in [31] by using Weighted Essentially Non-Oscillatary (WENO) [34] and Discontinuous Galerkin (DG) [11] methods for spatial discretization.
@@ 2 Conservation law and solution space
where $\bm{u}\in\mathbb{R}^{p}$ is the vector of conserved quantities, $\bm{f}(\bm{u})$ is the corresponding flux, together with some initial and boundary conditions. The solution that is physically correct is assumed to belong to an admissible set, denoted by $\mathcal{U}_{\textrm{$\operatorname{ad}$}}$. For example, in case of compressible flows, the density and pressure (or internal energy) must remain positive. In case of shallow water equations, the water depth must remain positive. In most of the models that are of interest, the admissible set is a convex subset of $\mathbb{R}^{p}$, and can be written as
@@
$\Omega_{e}=[x_{{e-\frac{1}{2}}},x_{{e+\frac{1}{2}}}]\qquad\textrm{and}\qquad\Delta x_{e}=x_{{e+\frac{1}{2}}}-x_{{e-\frac{1}{2}}}$
@@
which will be taken to be Gauss-Legendre (GL) or Gauss-Legendre-Lobatto (GLL) nodes, and will also be referred to as solution points. There are associated quadrature weights $w_{j}$ such that the quadrature rule is exact for polynomials of degree up to $12N+$ 1 for GL points and up to degree $12N-$ 1 for GLL points. Note that the nodes and weights we use are with respect to the interval $[0,1]$ whereas they are usually defined for the interval $[-1,+1]$. The solution inside an element is given by
@@ 3 Runge-Kutta FR
We first give a brief resume of RKFR scheme since our method has some similarities and borrows ideas from FR method. The RKFR scheme is based on an FR spatial discretization leading to a system of ODEs followed by application of an RK scheme to march forward in time. The key idea is to construct a continuous polynomial approximation of the flux which is then used in a collocation scheme to update the nodal solution values. At some time $t$, we have the piecewise polynomial solution defined inside each cell; the FR scheme can be described by the following steps.
@@
$\bm{f}_{h}(\xi,t)=\left[\bm{f}_{{e-\frac{1}{2}}}(t)-\bm{f}_{h}^{\delta}(0,t)\right]g_{L}(\xi)+\bm{f}_{h}^{\delta}(\xi,t)+\left[\bm{f}_{{e+\frac{1}{2}}}(t)-\bm{f}_{h}^{\delta}(1,t)\right]g_{R}(\xi)$ (4)
@@
We obtain the system of ODEs by collocating the PDE at the solution points
$\frac{\text{\text{{{d}}}}\bm{u}_{p}^{e}}{\text{\text{{{d}}}}t}(t)=-\frac{1}{\Delta x_{e}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\bm{f}_{h})_{\xi}}(\xi_{p},t),\qquad 0\leq p\leq N$
@@
Remark 2
The global flux approximation in (4) is globally continuous in 1-D. However, in higher dimensions, the flux is only made continuous along the normal direction to the element interfaces.
@@ 4 Multi-derivative Runge-Kutta FR scheme
Multiderivative Runge-Kutta [22] methods were initially developed to solve systems of ODEs like
@@
that use temporal derivatives of $\bm{L}$. In this work, we use the two stage fourth order multiderivative Runge-Kutta method from [20]. For the system of ODEs (5), the MDRK scheme of [20] to evolve from $t^{n}$ to $t^{n+1}$ is given by
@@
$\begin{split}\bm{F}&:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{f}_{t}}\left(\bm{u}^{n}\right)\\ \bm{F}^{\ast}&:=\bm{f}(\bm{u}^{n})+\frac{1}{6}\Delta t\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{f}_{t}}(\bm{u}^{n})+2{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{f}_{t}}(\bm{u}^{\ast})\right)\end{split}$ (8)
@@ 4.1 Conservation property
The computation of correct weak solutions for non-linear conservation laws in the presence of discontinuous solutions requires the use of conservative numerical schemes. The Lax-Wendroff theorem shows that if a consistent, conservative method converges, then the limit is a weak solution. The method (9) is also conservative though it is not directly apparent; to see this multiply (9) by the quadrature weights associated with the solution points and sum over all the points in the $e^{\text{th}}$ element,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{\bm{u}}_{e}^{*}:=}\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{\ast}$ $=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{2\Delta x_{e}}\sum_{p=0}^{N}w_{p}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\text{\text{{{d}}}}\bm{F}_{h}}{\text{\text{{{d}}}}\xi}}(\xi_{p})$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\overline{\bm{u}}_{e}^{n+1}:=}\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n+1}$ $=\sum_{p=0}^{N}w_{p}\bm{u}_{e,p}^{n}-\frac{\Delta t}{\Delta x_{e}}\sum_{p=0}^{N}w_{p}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\text{\text{{{d}}}}\bm{F}^{\ast}_{h}}{\text{\text{{{d}}}}\xi}}(\xi_{p})$ (10)
@@
where $\mathcal{U}_{\textrm{$\operatorname{ad}$}}$ is the admissibility set of the conservation law, and $\overline{\bm{u}}_{e}^{n+1}$ is element mean as given by (10).
@@ 4.3 Approximate Lax-Wendroff procedure
$\bm{F}:=\bm{f}\left(\bm{u}^{n}\right)+\frac{1}{4}\Delta t{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{f}_{t}}\left(\bm{u}^{n}\right)\approx\frac{1}{\Delta t/2}\int_{t^{n}}^{t^{n+1/2}}\bm{f}\left(\bm{u}\right)\text{\text{{{d}}}}t$ (14)
@@ 5.1 Stage 1
$\mathsf{\bm{F}}_{e}=a\mathsf{\bm{U}}_{e}\quad\text{where}\quad\mathsf{\bm{U}}_{e}=\mathsf{T}^{(1)}\mathsf{u}_{e}\quad\text{and}\quad\mathsf{T}^{(1)}=\mathsf{I}-\frac{\sigma}{4}\mathsf{D}$
@@ 5.2 Stage 2
and the flux derivative at the solution points is given by
@@ 6 Blending scheme
The MDRK scheme (9) gives a high (fourth) order method for smooth problems, but an important class of practical problems for hyperbolic conservation laws consist of non-smooth solutions with shocks and other discontinuities. In such situations, using a higher order method is bound to produce Gibbs oscillations, as stated in Godunov’s order barrier theorem [16]. The cure is to non-linearly add dissipation in regions where the solution is non-smooth, with methods like artificial viscosity, limiters and switching to a robust lower order scheme; the resultant scheme will be non-linear even for linear equations. In this work, we develop a subcell based blending scheme for MDRK similar to the one in [6] for the high order single stage Lax-Wendroff methods. The limiter is applied to each MDRK stage.
@@
where $\alpha_{e}\in[0,1]$ must be chosen based on the local smoothness indicator. If $\alpha_{e}=0$ then we obtain the high order MDRK scheme, while if $\alpha_{e}=1$ then the scheme becomes the low order scheme that is less oscillatory. In subsequent sections, we explain the details of the lower order scheme; for a description of the smoothness indicator, the reader is referred to [17, 6]. The lower order scheme will either be a first order finite volume scheme or a high resolution scheme based on MUSCL-Hancock idea. In either case, the common structure of the low order scheme can be explained as follows. In the first stage, the lower order residual $\mathsf{R}^{\ast,L}_{e}$ performs evolution from time $t^{n}$ to $t^{n+\frac{1}{2}}$ while, in the second stage, $\mathsf{R}^{L}_{e}$ performs evolution from $t^{n}$ to $t^{n+1}$. It may seem more natural to evolve from $t^{n+\frac{1}{2}}$ to $t^{n+1}$ in the next stage, but that will violate the conservation property because of the expression of second stage of MDRK (727).
@@
The MDRK scheme used in this work is fourth order accurate so we use polynomial degree $N=3$ in what follows, but write the procedure for general $N$. Let us subdivide each element $\Omega_{e}$ into $N+1$ subcells associated to the solution points $\{x^{e}_{p},p=0,1,\ldots,N\}$ of the MDRK scheme. Thus, we will have $N+2$ subfaces denoted by $\{x^{e}_{p+\frac{1}{2}},p=-1,0,\ldots,N\}$ with $x^{e}_{-\frac{1}{2}}=x_{{e-\frac{1}{2}}}$ and $x^{e}_{N+\frac{1}{2}}=x_{{e+\frac{1}{2}}}$. For maintaining a conservative scheme, the $p^{\text{th}}$ subcell is chosen so that
@@
where $w_{p}$ is the $p^{\text{th}}$ quadrature weight associated with the solution points. Figure 2 gives an illustration of the subcells for degree $N=3$ case. We explain the blending scheme for the second stage, where the lower order scheme performs evolution from $t^{n}$ to $t^{n+1}$. The low order scheme is obtained by updating the solution in each of the subcells by a finite volume scheme,
@@ 6.2 Higher order blending
7 Admissibility preservation
The Runge-Kutta FR schemes described in Section 3 using a Strong Stability Preserving (SSP) time discretization satisfy the admissibility in means property (Definition 2[46]. For any admissibility preserving in means FR scheme, the scaling limiter of [46] described in Section 7.2 can be used to obtain an admissibility preserving scheme (Definition 1). Since every stage of the multiderivative Runge-Kutta (MDRK) method uses a high order time-averaged flux (1415), the base MDRK scheme does not satisfy the admissibility preserving in means property. Thus, we propose to use the flux limiting procedure of [6] at every stage to choose the blended numerical fluxes (34) and ensure admissibility preservation in means for MDRK with blending scheme described in Section 6. The procedure for the MDRK scheme is briefly described in this section; for more details, the reader is referred to [6].
@@ 7.1 Flux limiter
The flux limiter to enforce admissibility will be applied in each stage so that the evolution of each stage is admissibility preserving in means. Thus, the admissibility enforcing flux limiting procedure of [6] applied to each stage of MDRK ensures admissibility in means of that stage and thus of the overall scheme.
The theoretical basis for flux limiting can be summarised in the following Theorem 1.
Theorem 1
Consider the MDRK scheme with blending (29) where low and high order schemes use the same numerical fluxes $\bm{F}_{e+\frac{1}{2}},\bm{F}^{\ast}_{e+\frac{1}{2}}$ (33) at every element interface. Then the following can be said about admissibility preserving in means property (Definition 2) of each stage of the MDRK scheme:
  1. 1.
    element means of both low and high order schemes are same and thus the blended scheme (29) is admissibility preserving in means if and only if the lower order scheme is admissibility preserving in means;
  2. 2.
    if the finite volume method using the lower order fluxes $\bm{f}_{e+\frac{1}{2}}$ (3536) as the interface flux is admissibility preserving, such as the first-order finite volume method or the MUSCL-Hancock scheme with CFL restrictions and slope correction from Theorem 3 of [6], and the blended numerical fluxes $\bm{F}_{e+\frac{1}{2}},\bm{F}^{\ast}_{e+\frac{1}{2}}$ are chosen to preserve the admissibility of lower-order updates at solution points adjacent to the interfaces, then the blending scheme (29) will preserve admissibility in means.
Proof.
By (10, 32), element means are the same for both low and high order schemes. Thus, admissibility in means of one implies the same for other, proving the first claim. For the second claim, note that our assumptions imply $(u_{j}^{e})^{L,n+1}$ given by (31) is in $\mathcal{U}_{\textrm{$\operatorname{ad}$}}$ for $0\leq j\leq N$, implying admissibility in means property of the lower order scheme by (32) and thus admissibility in means for the blended scheme. ∎
We now explain the procedure of ensuring that the update obtained by the lower order scheme in the second stage will be admissible. The application to the first stage is analogous. The lower order scheme is computed with a first order finite volume method or MUSCL-Hancock with slope correction from Theorem 3 of [6] so that admissibility is already ensured for inner solution points; i.e., we already have
$(\bm{u}^{e}_{j})^{L,n+1}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}},\quad 1\leq j\leq N-1$
The remaining admissibility constraints for the first ($j=0$) and last solution points ($j=N$) will be satisfied by appropriately choosing the inter-element flux $F_{{e+\frac{1}{2}}}$. The first step is to choose a candidate for $\bm{F}^{\ast}_{e+\frac{1}{2}}$ which is heuristically expected to give reasonable control on spurious oscillations, i.e.,
$\bm{F}^{\ast}_{{e+\frac{1}{2}}}=(1-\alpha_{{e+\frac{1}{2}}})\bm{F}^{\text{HO}\ast}_{{e+\frac{1}{2}}}+\alpha_{{e+\frac{1}{2}}}\bm{f}_{{e+\frac{1}{2}}},\quad\alpha_{{e+\frac{1}{2}}}=\frac{\alpha_{e}+\alpha_{e+1}}{2}$
where $\bm{f}_{{e+\frac{1}{2}}}$ is the lower order flux at the face ${e+\frac{1}{2}}$ shared between FR elements and subcells (35, 36), and $\alpha_{e}$ is the blending coefficient (29) based on element-wise smoothness indicator from [17] which is also described in Section 3.3 of [6].
The next step is to correct $\bm{F}^{\ast}_{{e+\frac{1}{2}}}$ to enforce the admissibility constraints. The guiding principle of our approach is to perform the correction within the face loops, minimizing storage requirements and additional memory reads. The lower order updates in subcells neighbouring the ${e+\frac{1}{2}}$ face with the candidate flux are
$\begin{split}\tilde{\bm{u}}_{0}^{n+1}&=(\bm{u}^{e+1}_{0})^{n}-\frac{\Delta t}{w_{0}\Delta x_{e+1}}(\bm{f}^{e+1}_{\frac{1}{2}}-\bm{F}^{\ast}_{{e+\frac{1}{2}}})\\ \tilde{\bm{u}}_{N}^{n+1}&=(\bm{u}^{e}_{N})^{n}-\frac{\Delta t}{w_{N}\Delta x_{e}}(\bm{F}^{\ast}_{{e+\frac{1}{2}}}-\bm{f}^{e}_{N-\frac{1}{2}})\end{split}$ (37)
In order to correct the interface flux, we will again use the fact that first order finite volume method and MUSCL-Hancock with slope corrections from [6] preserve admissibility, i.e.,
[…]
Let $\{p_{k},1\leq 1\leq K\}$ be the admissibility constraints (2) of the conservation law. The numerical flux is corrected by iterating over the admissibility constraints as follows
$\bm{F}^{\ast}_{{e+\frac{1}{2}}}\leftarrow(1-\alpha_{{e+\frac{1}{2}}})\bm{F}^{\text{HO}\ast}_{{e+\frac{1}{2}}}+\alpha_{{e+\frac{1}{2}}}\bm{f}_{{e+\frac{1}{2}}}$
for $k$=1:$K$ do
[…]
     $\theta\leftarrow\min\left(\min_{j=0,N}\left|\frac{\epsilon_{j}-p_{k}(\tilde{\bm{u}}_{j}^{\text{low},n+1})}{p_{k}(\tilde{\bm{u}}_{j}^{n+1})-p_{k}(\tilde{\bm{u}}_{j}^{\text{low},n+1})}\right|,1\right)$
     $\bm{F}^{\ast}_{{e+\frac{1}{2}}}\leftarrow\theta\bm{F}^{\ast}_{{e+\frac{1}{2}}}+(1-\theta)\bm{f}_{{e+\frac{1}{2}}}$
     $\tilde{\bm{u}}_{0}^{n+1}\leftarrow(\bm{u}^{e+1}_{0})^{n}-\frac{\Delta t}{w_{0}\Delta x_{e+1}}(\bm{f}^{e+1}_{\frac{1}{2}}-\bm{F}^{\ast}_{e+\frac{1}{2}})$
     $\tilde{\bm{u}}_{N}^{n+1}\leftarrow(\bm{u}^{e}_{N})^{n}-\frac{\Delta t}{w_{N}\Delta x_{e}}(\bm{F}^{\ast}_{e+\frac{1}{2}}-\bm{f}^{e}_{N-\frac{1}{2}})$
end for
By concavity of $p_{k}$, after the $k^{\text{th}}$ iteration, the updates computed using flux $\bm{F}^{\ast}_{e+\frac{1}{2}}$ will satisfy
[…]
7.2 Scaling limiter
The flux limiter described in the previous section gives us an admissibility preserving in means MDRK scheme. In this section, we review the scaling limiter of [46] to use the admissibility in means property (Definition 2) to obtain an admissibility preserving scheme (Definition 1). Consider the solution $\bm{u}_{h}^{n}$ at the current time time level $n$. Within each element, $\bm{u}_{h}^{n}\in\mathbb{P}_{N}$ and since the scheme is admissibility preserving in means, we assume $\overline{\bm{u}}^{n}_{e}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}$ for each element $e$. We will iteratively correct all admissibility constraints $\left\{p_{k}\right\}_{k=1}^{K}$ (2). For each constraint $p_{k}$, we find $\theta_{k}\in[0,1]$ such that $p_{k}\left((1-\theta_{k})\overline{\bm{u}}^{n}_{e}+\theta_{k}\bm{u}_{h}^{n}\right)>0$ at the $N+1$ solution points and replace the polynomial $\bm{u}_{h}^{n}$ with $(1-\theta)\overline{\bm{u}}^{n}_{e}+\theta\bm{u}_{h}^{n}$ where $\theta=\min_{k}\theta_{k}$. This correction procedure is performed for all admissibility constraints as follows.   $\theta=1$ for $k$=1:$K$ do      $\epsilon_{k}\leftarrow\frac{1}{10}p_{k}(\overline{\bm{u}}_{e}^{n})$      $\theta_{k}=\min\left(\min_{0\leq j\leq N}\left|\frac{\epsilon_{p}-p_{k}(\overline{\bm{u}}^{n}_{e})}{p_{k}(\bm{u}_{e,j}^{n})-p_{k}(\overline{\bm{u}}^{n}_{e})}\right|,1\right)$      $\theta\leftarrow\min(\theta_{k},\theta)$ end for   By the same arguments as (38), the procedure will ensure $\bm{u}_{e,j}^{n}\in\mathcal{U}_{\textrm{$\operatorname{ad}$}}$ for all $e,j$ by the concavity of the admissibility constraints $\{p_{k}\}$ (2). Thus, by Definition 2, the element means at the next time level will also be admissible. Overall, by performing this procedure at every time level, we have an admissibility preserving scheme in the sense of Definition 1.